# Auditing Student–AI Collaboration: A Case Study of Online Graduate CS Students

### v2

@@ Abstract.
As generative AI becomes embedded in higher education, it increasingly shapes how students complete academic tasks. While these systems offer efficiency and support, concerns persist regarding over-automation, diminished student agency, and the potential for unreliable or hallucinated outputs. This study conducts a mixed-methods audit of student–AI collaboration preferences by examining the alignment between current AI capabilities and students’ desired levels of automation in academic work. Using two sequential and complementary surveys, we capture students’ perceived benefits, risks, and preferred boundaries when using AI. The first survey employs an existing task-based framework to assess preferences for and actual usage of AI across 12 academic tasks, alongside primary concerns and reasons for use. The second survey, informed by the first, explores how AI systems could be designed to address these concerns through open-ended questions. This study aims to identify gaps between existing AI affordances and students’ normative expectations of collaboration, informing the development of more effective and trustworthy AI systems for education.
AI in education, human–AI collaboration, student agency, automation preferences, generative AI, academic integrity, HCAI
copyright: acmlicensedjournalyear: 2026doi: 10.1145/XXXXXXX.XXXXXXXconference: Conference Full Name; June 03–05, 2026; Woodstock, NYbooktitle: Conference Full Name (Conference ’26), June 03–05, 2026, Woodstock, NYisbn: 978-1-4503-XXXX-X/26/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Generative artificial intelligence (GenAI) tools such as ChatGPT, Claude, and Gemini are rapidly becoming embedded in higher education, reshaping how students complete academic tasks (Kasneci et al., 2023; Chan and Hu, 2024; Sukumar et al., 2025). These systems promise substantial gains in efficiency, personalization, and support, yet they also introduce significant risks related to over-automation, reduced student agency, and vulnerability to unreliable or hallucinated outputs (Shneiderman, 2020; Ji et al., 2023; Athaluri et al., 2023). As students increasingly delegate cognitive and procedural work to AI, it becomes essential to understand not only how they use these tools, but also how they perceive the trade-offs between automation and autonomy, and what design features would make AI systems more trustworthy in educational contexts (Khosravi et al., 2022; Amershi et al., 2019).
@@
Recent research highlights the dual-edged nature of GenAI in education. Systematic reviews indicate that these tools can enhance writing quality, support personalized learning, and increase productivity, while simultaneously raising concerns about academic integrity, bias, and equitable access (Sukumar et al., 2025; Li et al., 2024; Camacho-Zúñiga et al., 2025). Empirical studies further suggest that the impact of AI depends heavily on how it is integrated pedagogically (Lee and Moore, 2024; Maizatul Akmar et al., 2025). However, a critical gap remains in understanding how students themselves navigate collaboration with AI across diverse academic tasks, particularly in graduate-level computer science education where technical accuracy and critical reasoning are paramount (Dwivedi et al., 2023; Sullivan et al., 2023).
@@
This study addresses that gap by conducting a mixed-methods audit of student–AI collaboration among graduate students in the Georgia Tech Online Master of Science in Computer Science (OMSCS) program. Grounded in a Human-Centered AI (HCAI) perspective (Shneiderman, 2020, 2022) and informed by the Human Agency Scale (HAS) (Shao et al., 2025), we investigate two core dimensions of this collaboration. First, we measure students’ preferred levels of automation across twelve common academic tasks—spanning reading, writing, coding, studying, collaboration, and assessment—and compare these preferences to their actual AI usage (Shao et al., 2025, 2024b). This allows us to map tasks into four alignment zones: Green Light (high desire, high use), R&D Opportunity (high desire, low use), Low Priority (low desire, low use), and Red Light (low desire, high use). Second, through qualitative analysis of open-ended responses, we identify the system-level features—such as transparency, confidence indicators, explainability, hallucination warnings, and pedagogical alignment—that students believe would enhance the trustworthiness of AI in education (Amershi et al., 2019; Khosravi et al., 2022, 2024).
By integrating quantitative survey data with qualitative insights, this study provides a detailed, person-centered view of how graduate CS students perceive and negotiate the opportunities and risks of AI collaboration. Our findings contribute to the growing discourse on bidirectional human–AI alignment in education (Shen et al., 2024), offering evidence-based recommendations for educators, designers, and policymakers seeking to foster responsible, human-centered AI adoption in higher learning environments (Camacho-Zúñiga et al., 2025; Khosravi et al., 2024).
In this paper, we investigate the following research questions:
@@
To answer these questions, we administered a two-part survey to over 50 OMSCS students. Part 1 employed a task-based adaptation of the Human Agency Scale (Shao et al., 2025) to quantify automation preferences and actual usage, as well as students’ primary reasons for and concerns about AI use across tasks. Part 2 used an open-ended prompt to elicit qualitative feedback on features that would address these concerns in educational AI (Braun and Clarke, 2006). Our analysis integrates these data to produce an automation alignment map and a thematic framework for trustworthy AI design.
Our findings indicate that students strongly desire automation for repetitive, time-consuming, or accuracy-focused tasks, but remain cautious about delegating activities that involve higher-order reasoning, creativity, or technical nuance (Cotton et al., 2023; Halaweh, 2023). Notably, no tasks fell into the Red Light Zone, suggesting that current usage does not exceed students’ comfort levels. Students also consistently called for greater transparency, confidence indicators, source citations, and safeguards against hallucinations—preferences that echo prior calls for explainable and controllable AI systems (Amershi et al., 2019; Khosravi et al., 2022, 2024).
This research makes three main contributions:
  • Empirical: We provide a detailed, mixed-methods audit of student–AI collaboration in a real-world graduate CS context, capturing both behavioral patterns and perceived trust factors.
  • Methodological: We demonstrate the utility of the Human Agency Scale (Shao et al., 2025) and alignment zone mapping for evaluating automation acceptance in education.
  • Practical: We offer evidence-based design implications and policy recommendations for creating more transparent, reliable, and human-centered AI tools in higher education (Amershi et al., 2019; Shneiderman, 2022).
@@ 2. Related Work
Research on artificial intelligence in education has expanded rapidly with the emergence of large-scale generative language models. Prior work spans educational research, learning sciences, human–computer interaction, and socio-technical studies of automation, addressing both opportunities and risks of AI-supported learning. This section synthesizes three strands most relevant to our study: (1) empirical and ethical research on generative AI in higher education, (2) Human-Centered AI (HCAI) as a design framework for trustworthy educational technologies, and (3) approaches to measuring human agency and automation preferences, emphasizing the Human Agency Scale (HAS) as a recent operationalization for quantifying desired human involvement.
@@ 2.1. Generative AI in Higher Education
At the same time, substantial concerns have been raised about over-reliance on AI, hallucinated or biased outputs, and the potential erosion of critical thinking and disciplinary epistemic skills (Bastani et al., 2024; Selwyn, 2019). Empirical work suggests that when students delegate substantive intellectual work—such as complex problem solving or argumentation—to AI systems without appropriate scaffolding, learning outcomes can suffer (Bastani et al., 2024). These risks are particularly salient in technically demanding disciplines such as computer science, where correctness, abstraction, and conceptual understanding are central to education.
@@
In addition to learning outcomes, generative AI raises fundamental challenges for academic integrity, authorship, and equity. Scholars argue that traditional assessment frameworks are increasingly misaligned with AI-enabled workflows, driving the need for new assessment designs and integrity policies (Perkins, 2023b). Disparities in access to AI tools and differences in AI literacy further risk exacerbating educational inequalities, reinforcing calls for responsible, inclusive integration of AI in education (Holmes et al., 2022). Together, this literature underscores the need for approaches that harness efficiency gains while preserving autonomy, integrity, and equitable participation.
@@ 2.2. Human-Centered AI in Educational Contexts
A concept in HCAI is bidirectional alignment: AI systems should be designed to reflect human values and pedagogical aims, while users must develop the critical capacity to engage with AI systems responsibly (Shen et al., 2024). In educational settings, this mutual adaptation matters because student–AI interactions can shape epistemic beliefs, study approaches, and perceptions of authorship and ownership of work. Our work adopts this perspective by conceptualizing student–AI collaboration as an ongoing partnership requiring thoughtful system design and active human engagement.
@@ 2.3. Human Agency and Automation Preferences
As AI systems become active collaborators in human workflows, researchers have sought to quantify how users perceive and negotiate their agency under automation. A recent contribution to this literature is the **Human Agency Scale (HAS)** introduced by Shao et al., which provides a five-level scale (H1–H5) to quantify the degree of human involvement preferred in task performance when working with AI agents (Shao et al., 2025). HAS moves beyond binary automate-or-not frameworks by offering a shared language for examining desired human control versus automation or augmentation across diverse tasks and occupations.
The HAS differentiates between complete human control (higher levels) and higher degrees of automation (lower levels), allowing for nuanced analyses of user preferences in human–AI work contexts. Validation and application of HAS indicate that agency perceptions vary systematically with task criticality, trust, perceived AI competence, and potential consequences of error, making the scale well-suited to assess collaborative human–AI workflows. Related research on algorithm aversion and resistance to automation shows that users are often reluctant to delegate control in domains tied to expertise, identity, or accountability (Dietvorst et al., 2015; Longoni et al., 2019).
Within educational research, HAS-aligned findings suggest that students are typically comfortable delegating repetitive or mechanical tasks — such as formatting, grammar correction, or simple content summarization — while preferring to retain control over tasks that involve creativity, reasoning, or deep disciplinary judgment (Liu et al., 2024; Bastani et al., 2024). These preferences are shaped by trust, perceived risk, and confidence in one’s ability to evaluate AI outputs — critical factors that influence appropriate reliance on automated systems.
Our study adapts the HAS framework to graduate computer science education, applying it at a task-specific level to compare desired human agency, actual AI use, and confidence in verifying outputs. This approach enables a fine-grained audit of student–AI collaboration and provides empirical grounding for human-centered design recommendations.
2.4. Research Gap
Taken together, prior work establishes the promise of generative AI for education, the pedagogical and ethical risks it poses, and the importance of agency-preserving, human-centered design. However, much existing research focuses either on macro-level outcomes (e.g., policy debates, institutional assessment practices) or on system-centric evaluations of AI performance. There is limited attention to how students themselves experience and negotiate collaboration with AI across concrete academic tasks, especially in disciplines with strong technical demands.
Few studies integrate quantitative measures of automation preference with qualitative insights into trust, transparency, and desired system features. By conducting a mixed-methods audit of student–AI collaboration grounded in HCAI principles and operationalized through the Human Agency Scale, our work addresses this gap and contributes a person-centered perspective to the design, adoption, and governance of AI in higher education.
@@ 3. Methodology
Table 1. 12 academic tasks used in the study.
@@
3.1. 12 Academic Tasks
To systematically evaluate student preferences for AI collaboration, we selected twelve representative academic tasks that span the core cognitive and procedural activities common in graduate-level computer science education (Swiecki et al., 2022). These tasks were drawn from six functional categories: (1) Reading and Note-Making, (2) Writing and Revision, (3) Coding and Problem Solving, (4) Studying and Metacognition, (5) Collaboration and Communication, and (6) Assessment and Feedback. This categorization reflects the multifaceted nature of graduate learning, where students engage not only in content acquisition but also in synthesis, creation, and critical evaluation (Kasneci et al., 2023).
Each task was chosen based on its relevance to both conceptual and practical skill development. For instance, tasks such as debugging code (T5) and solving quantitative problems step-by-step (T10) represent core technical competencies, while activities like summarizing research (T1) and providing feedback on writing (T12) emphasize critical engagement and communication skills (Shao et al., 2024a). By including tasks that range from routine (e.g., formatting citations) to complex and open-ended (e.g., brainstorming essay structures), we aimed to capture a spectrum of automation preferences and perceived risks.
The tasks were designed to vary along several dimensions relevant to AI collaboration:
Cognitive Demand: From low-level, rule-based tasks (T2, T4) to high-level, interpretive tasks (T3, T12) (Swiecki et al., 2022).
Structured vs. Open-Ended: Well-defined tasks with clear correctness criteria (T5, T10) versus tasks requiring creativity or subjective judgment (T3, T6) (Mollick and Mollick, 2023).
Frequency and Time Cost: Common, time-consuming activities (T1, T11) versus occasional but high-stakes tasks (T9, T12).
Potential for AI Error: Tasks where AI hallucinations or inaccuracies could have meaningful consequences (T1, T5, T10) (Ji et al., 2023; Athaluri et al., 2023).
This variation allows us to examine how task characteristics shape students’ willingness to delegate work to AI, their confidence in verifying outputs, and their primary concerns regarding reliability and learning integrity (Cotton et al., 2023; Halaweh, 2023).
@@
3.2. Desired Automation Likert Questions
@@
Figure 1. Illustration of the Likert-scale question used to measure students’ desired level of AI automation for academic tasks.
@@
3.3. Actual Usage Likert Question
Likert Question for Observed AI Usage Question. For each academic task listed below, please indicate how capable current AI systems are based on your own experience, reflected by how you actually use AI for that task. Scale. 1: I do not use AI for this task at all 2: I use AI to provide minor or partial help 3: I use AI for this task moderately 4: AI handles most of the task with human oversight 5: I rely entirely on AI to automate this task
Figure 2. Illustration of the Likert-scale question used to capture students’ self-reported observed AI usage across academic tasks.
@@ 3.3. Actual Usage Likert Question
While self-reported usage does not substitute for expert evaluation, it provides a meaningful proxy for how capable AI systems are perceived to be in real educational settings, particularly in contexts where students make day-to-day decisions about whether and how much to rely on AI. Maintaining the same automation continuum as the desired automation measure allows us to directly compare normative expectations (how AI should be used) with descriptive practices (how AI is actually used).
@@ 3.4. Four-Zone Classification of AI Automation
Automation “Red Light” Zone. The Red Light zone includes tasks that participants believe should not be automated, even if AI systems are technically capable of performing them. This zone reflects strong human resistance to AI involvement, often due to concerns related to trust, accountability, ethical implications, or academic integrity (Cotton et al., 2023; Perkins, 2023a). Tasks in this zone signal boundaries where AI use is perceived as inappropriate despite potential performance gains.
@@
In our study, participants’ Likert-scale responses regarding desired levels of AI automation and confidence in AI-generated outputs were jointly considered to map academic tasks onto these four zones. This classification enabled a systematic comparison of where students perceive AI use as appropriate, inappropriate, promising but underdeveloped, or low priority within the context of AI-assisted coursework.
@@ 3.5. Primary Concerns and Reasons for AI Usage across Academic Tasks
To understand what shapes students’ adoption of AI across different academic tasks, we asked participants to report both their primary concerns when using AI and their reasons for relying on AI assistance (Dwivedi et al., 2023; Sullivan et al., 2023). Concerns included risks such as inaccurate or misleading outputs, hallucinations, academic misconduct, and reduced critical thinking (Cotton et al., 2023; Halaweh, 2023; Imran and Almusharraf, 2023), while reasons for use captured perceived benefits such as saving time, reducing cognitive load while managing conceptually complex task and improving accuracy or polishing academic outputs (Kasneci et al., 2023).
Analyzing these responses at the task level allows us to examine how motivations for AI use coexist with, and are constrained by, students’ concerns. While students often report using AI to support efficiency and reduce effort in repetitive or demanding tasks, concerns about hallucinations and reliability remain salient, particularly for tasks involving complex reasoning or high academic stakes (Ji et al., 2023; Athaluri et al., 2023; Weiser, 2023). This dual perspective highlights that students’ AI use is not driven solely by convenience, but reflects ongoing trade-offs between perceived benefits and epistemic risks (Bender et al., 2021).
These findings provide important context for our subsequent analysis, motivating a closer examination of how students experience AI hallucinations in practice and what system-level design features they expect to better address these concerns.
@@ 3.6. Design Expectations for Addressing AI Concerns
In addition to structured Likert-scale items, we included open-ended questions to elicit students’ perspectives on how AI systems should be designed to address their concerns in educational contexts (Braun and Clarke, 2006). These questions focused on design features and interaction mechanisms that could improve trustworthiness, support error detection, and mitigate risks such as hallucination, without constraining responses to predefined options (Perkins, 2023a).
The use of open-ended questions allows participants to express design expectations in their own terms, capturing aspects of AI system behavior and interface design that may not be fully anticipated by closed-form survey items (Braun and Clarke, 2006). We treat these responses as qualitative inputs for identifying recurring design considerations and informing subsequent analysis of user-centered approaches to AI alignment in education (Swiecki et al., 2022).
@@ 4.1. Participants
Participants were graduate students enrolled in the Online Master of Science in Computer Science (OMSCS) program at the Georgia Institute of Technology, specifically students taking CS6460: Educational Technology. Participation in the survey was voluntary and counted toward a course participation grade.
To ensure the relevance of responses to AI-assisted coursework, we filtered out participants who reported rarely using or not using AI tools in their academic work. In total, $N=57$ students completed the survey. After filtering, $N=44$ participants who reported at least occasional use of AI systems for coursework were retained for analysis. The final sample therefore consists of students with meaningful experience interacting with AI in educational contexts, allowing us to focus our analysis on substantive patterns of AI use.

### v3

@@ Abstract.
As generative AI becomes embedded in higher education, it increasingly shapes how students complete academic tasks. While these systems offer efficiency and support, concerns persist regarding over-automation, diminished student agency, and the potential for unreliable or hallucinated outputs. This study conducts a mixed-methods audit of student–AI collaboration preferences by examining the alignment between current AI capabilities and students’ desired levels of automation in academic work. Using two sequential and complementary surveys, we capture students’ perceived benefits, risks, and preferred boundaries when using AI. The first survey employs an existing auditing framework to assess preferences for and actual usage of AI across 12 tasks in educational contexts, alongside primary concerns and reasons for use. The second survey, informed by the first, explores how AI systems could be designed to address these concerns through open-ended questions. This study aims to identify gaps between existing AI affordances and students’ normative expectations of collaboration, informing the development of more effective and trustworthy AI systems for education.
AI in education, human–AI collaboration, student agency, automation preferences, generative AI
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Proceedings of the Bilaign Workshop at CHI 2026; April 2026; Barcelona, Spainbooktitle: Proceedings of the BiAlign Workshop at CHI 2026isbn: 978-1-4503-XXXX-X/26/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Generative artificial intelligence (GenAI) tools such as ChatGPT, Claude, and Gemini are rapidly becoming embedded in higher education, reshaping how students complete academic tasks (Kasneci et al., 2023; Chan and Hu, 2025; Sukumar et al., 2025). These systems promise substantial gains in efficiency, personalization, and support, yet they also introduce significant risks related to over-automation, reduced student agency, and vulnerability to unreliable or hallucinated outputs (Shneiderman, 2020; Ji et al., 2023; Athaluri et al., 2023). As students increasingly delegate cognitive and procedural work to AI, it becomes essential to understand not only how they use these tools, but also how they perceive the trade-offs between automation and autonomy, and what design features would make AI systems more trustworthy in educational contexts (Khosravi et al., 2022; Amershi et al., 2019).
@@
Recent research highlights the dual-edged nature of GenAI in education. Systematic reviews indicate that these tools can enhance writing quality, support personalized learning, and increase productivity, while simultaneously raising concerns about academic integrity, bias, and equitable access (Sukumar et al., 2025; Li et al., 2024; Camacho-Zúñiga et al., 2025). Empirical studies further suggest that the impact of AI depends heavily on how it is integrated pedagogically (Lee and Moore, 2024; Maizatul Akmar et al., 2025). However, a critical gap remains in understanding how students themselves navigate collaboration with AI across diverse tasks in educational setting.
@@
This study addresses that gap by conducting a mixed-methods audit of student–AI collaboration among graduate students in the Georgia Tech Online Master of Science in Computer Science (OMSCS) program. Grounded in a Human-Centered AI (HCAI) perspective (Shneiderman, 2020, 2022) and informed by the Human Agency Scale (HAS) (Shao et al., 2025), we investigate two core dimensions of this collaboration. First, we measure students’ preferred levels of automation across twelve common tasks in educational contexts and compare these preferences to their actual AI usage (Shao et al., 2025, 2024b). This allows us to map tasks into four alignment zones: Green Light (high desire, high use), R&D Opportunity (high desire, low use), Low Priority (low desire, low use), and Red Light (low desire, high use). Second, through qualitative analysis of open-ended responses, we identify the system-level features that students believe would enhance the trustworthiness of AI in education (Amershi et al., 2019; Khosravi et al., 2022, 2024).
In this paper, we investigate the following questions:
@@ 2.1. Generative AI in Higher Education
At the same time, substantial concerns have been raised about over-reliance on AI, hallucinated or biased outputs, and the potential erosion of critical thinking and disciplinary epistemic skills (Bastani et al., 2024; Selwyn, 2019). Empirical work suggests that when students delegate substantive intellectual work—such as complex problem solving or argumentation—to AI systems without appropriate scaffolding, learning outcomes can suffer (Bastani et al., 2024).
@@ 2.3. Human Agency and Automation Preferences
As AI systems become active collaborators in human workflows, researchers have sought to quantify how users perceive and negotiate their agency under automation. A recent contribution to this literature is the Human Agency Scale (HAS) introduced by Shao et al., which provides a five-level scale (H1–H5) to quantify the degree of human involvement preferred in task performance when working with AI agents (Shao et al., 2025). HAS moves beyond binary automate-or-not frameworks by offering a shared language for examining desired human control versus automation or augmentation across diverse tasks and occupations.
Our study adapts the HAS framework to education domain, applying it at a task-specific level to compare desired human agency and actual AI use. This approach enables a fine-grained audit of student–AI collaboration and provides empirical grounding for human-centered design recommendations.
@@ 3. Methodology
Table 1. 12 tasks in educational contexts used in the study.
@@
3.1. 12 Tasks in Educational Context
To systematically evaluate student preferences for AI collaboration, we selected twelve tasks in educational setting. Each task was chosen based on its relevance to both conceptual and practical skill development. For instance, tasks such as debugging code (T5) and solving quantitative problems step-by-step (T10) represent core technical competencies, while activities like summarizing research (T1) and providing feedback on writing (T12) emphasize critical engagement and communication skills (Shao et al., 2024a). By including tasks that range from routine (e.g., formatting citations) to complex and open-ended (e.g., brainstorming essay structures), we aimed to capture a spectrum of automation preferences and perceived risks.
@@
3.2. Desired vs. Actual Automation Likert Questions
@@
Figure 1. Desired level of AI automation.
Likert Question for Collecting Observed AI Usage Question. For each academic task listed below, please indicate how capable current AI systems are based on your own experience, reflected by how you actually use AI for that task. Scale. 1: I do not use AI for this task at all 2: I use AI to provide minor or partial help 3: I use AI for this task moderately 4: I use AI to handle most of the task with oversight 5: I rely entirely on AI to automate this task
Figure 2. Actual AI usage Level.
@@ 3.3. Four-Zone Classification of AI Automation
Automation “Red Light” Zone. The Red Light zone includes tasks that participants believe should not be automated, even if AI systems are technically capable of performing them. This zone reflects strong human resistance to AI involvement, often due to concerns related to trust, accountability, ethical implications, or academic integrity (Cotton et al., 2023; Perkins, 2023). Tasks in this zone signal boundaries where AI use is perceived as inappropriate despite potential performance gains.
@@ 3.4. Primary Concerns and Reasons for AI Usage across Academic Tasks
To understand what shapes students’ adoption of AI across different academic tasks, we asked participants to report both their primary concerns when using AI and their reasons for relying on AI assistance (Dwivedi et al., 2023; Sullivan et al., 2023). Concerns included risks such as inaccurate or misleading outputs, hallucinations, academic misconduct, and reduced critical thinking (Cotton et al., 2023; Halaweh, 2023; Imran and Almusharraf, 2023), while reasons for use captured perceived benefits such as saving time, reducing cognitive load while managing conceptually complex task and improving accuracy or polishing academic outputs (Kasneci et al., 2023), as shown in Table 2.
@@ 3.5. Design Expectations for Addressing AI Concerns
In addition to structured Likert-scale items, we included open-ended question to elicit students’ perspectives on how AI systems should be designed to address their greatest concern in educational contexts (Braun and Clarke, 2006). The use of open-ended questions allows participants to express design expectations in their own terms, capturing aspects of AI system behavior and interface design that may not be fully anticipated by closed-form survey items (Braun and Clarke, 2006). We treat these responses as qualitative inputs for identifying recurring design considerations and informing subsequent analysis of user-centered approaches to AI alignment in education (Swiecki et al., 2022).
@@ 4.1. Participants
Participants were graduate students enrolled in the Online Master of Science in Computer Science (OMSCS) program at the Georgia Institute of Technology. Participation in the survey was voluntary. To ensure the relevance of responses to AI-assisted coursework, we filtered out participants who reported rarely using or not using AI tools in their academic work. In total, $N=57$ students completed the survey. After filtering, $N=44$ participants who reported at least occasional use of AI systems for coursework were retained for analysis. The final sample therefore consists of students with meaningful experience interacting with AI in educational contexts, allowing us to focus our analysis on substantive patterns of AI use.